OpenStudy (anonymous):

A customer has a debit card that gives monthly cash back bonuses. If the customer spends from $10 to $499.99 he earns 1% cash back. If he spends $500 to $999.99 he earns 2% cash back. For monthly purchases of $1000 and more, he earns a 3% cash back bones. His monthly purchases for the past 6 months are as follows: $345, $609. $280, 1020 and $45. How much did he earn in cash bonuses over the past six months?

OpenStudy (anonymous):

The first month he spends $345. Since that is less than $499.99, he gets 1% cash back on $345-$10. So (345-10)*.01=3.35 So he earns $3.35 cash back for the first month. The second month he spends $609. So the first $499.99 spent earns 1% cash back. And the addition amount he spends earns 2% cash back. So (609-499.99).02 + (499.99-10)*.01 = 2.18 +4.90 = 7.08 So he earns $7.08 for the second month. See if you can finish the problem now.